OpenStudy (shall29):

For the second order reaction: A --> products the plot of what term versus time would yield a straight line?

OpenStudy (shall29):

@Photon336 @Prathamesh_M @Preetha @sweetburger

OpenStudy (shall29):

[A] 1/[A] ln[A]/[A] ln[A]

OpenStudy (prathamesh_m):

try solving the differential equation: \[-\frac{ d[A] }{ dt }=[A]^{2}\]

OpenStudy (prathamesh_m):

u will eventually end up with a relation between [A], i.e., concentration of A, and time

OpenStudy (shall29):

Okay I think the answer is either C or D.

OpenStudy (shall29):

What do you think?

OpenStudy (shall29):

@Prathamesh_M

OpenStudy (prathamesh_m):

What equation did u end up with finally?

OpenStudy (prathamesh_m):

I think u should try again i got B

OpenStudy (prathamesh_m):

\[-\frac{ d[A] }{ [A]^{2} }=dt\]

OpenStudy (prathamesh_m):

Now integrate on both sides applying appropriate limits

OpenStudy (shall29):

Okay, I got B now. @Photon336 what do you think, is that right?

OpenStudy (photon336):

well. first a second order rate law would probably look like this \[[A] \rightarrow products \] since we've got only reactant A becoming a product. we automatically know that the reaction must be second order with respect to A. \[r = k[A]^{2}\] but I believe that the rate is like how fast your reaction rate is. but it also means that [A] is disappearing at some rate. like we know the rate constant depends on time right? \[[A_{0}]^{2}k~dt = -d[A_{0}]\] so like what we're trying to say here is that I think that the rate of formation of products is equal to the rate we're losing A \[\frac{ -d[A_{0}] }{ [A_{0}]^{2} } = k~dt \] \[\int\limits \frac{ -d[A_{0}] }{ [A_{0}]^{2} } = \int\limits k~dt \] \[-\int\limits \frac{ 1 }{ u^{2} } = -1*-1 \frac{ 1 }{ u }~du => \frac{ 1 }{ [A_{0}] }\] hmm yeah this is what i'm getting. \[\frac{ 1 }{ [A_{0}] } = k~t\]

OpenStudy (photon336):

@shall29 I forgot to include something i put in the initial concentration but didn't take into consideration the final one. \[\int\limits_{[A]_{0}}^{[A]} \frac{ 1 }{ [A]^{2} } = \frac{ 1 }{ [A] }-\frac{ 1 }{ [A_{0}] }\] \[\frac{ 1 }{ [A] }-\frac{ 1 }{ [A_{0}] } = kt\] \[kt+\frac{ 1 }{ [A]_{0} } = \frac{ 1 }{ [A] }\]

OpenStudy (shall29):

Sorry about that. Okay I understand that now

OpenStudy (shall29):

So the answer is B, yes?

OpenStudy (prathamesh_m):

yes its 1/[A]